# From Reasoning LLMs to BERT: A Two-Stage Distillation Framework for Search Relevance

### v2

@@ 4.1. Datasets and Evaluation
We employ three metrics to evaluate model performance.
  • Accuracy is the fraction of correct predictions over the total number of samples, defined as:
    (9) $\text{Accuracy}(y,\hat{y})=\frac{1}{n_{\text{samples}}}\sum_{i=1}^{n_{\text{samples}}}\mathbf{1}(\hat{y}_{i}=y_{i})$
    where $\hat{y}_{i}$ is the prediction for the $i$-th sample, $y_{i}$ is the true value, and $\mathbf{1}(\cdot)$ is the indicator function.
  • Macro F1 computes the unweighted average of the F1-scores for each class, treating all classes equally. The base F1-score is calculated as $F_{1}=2\cdot\frac{P\cdot R}{P+R}$.
  • Weighted F1 computes a weighted average of the F1-scores, where each class’s score is weighted by its support. It is defined as $F_{1,\text{weighted}}=\sum_{c=1}^{C}w_{c}\cdot F_{1,c}$.
@@ 4.2. Implementation Details
We selected qwen3-8b (Yang et al., 2025b) as the base model. During the CPT stage, training was conducted with a batch size of 32 and a learning rate of 1e-5, which was managed by a cosine scheduler with a warm-up ratio of 0.1. The AdamW optimizer was employed. The model was trained for 2 epochs with a maximum input length of 2048 tokens, and the final loss converged to approximately 0.6. The hyperparameters for SFT were identical to those used in CPT, with the exception of the batch size were 16 and the number of epochs was set to 3, respectively. The reward model was trained using the same hyperparameters as the SFT stage. For the subsequent reinforcement learning (RL) stage, we employed VERL framework (Sheng et al., 2025) . During training, we sampled 16 responses per prompt, using a prompt batch size of 64 (mini-batch size of 32) and use AdamW optimizer with learning rate setting to 6e-7. To prevent GPU out-of-memory (OOM) errors, we deployed the reward model using vLLM (Kwon et al., 2023) and invoked it via API calls during the RL training process.
@@
For the distillation experiments, we employed a 6-layer Chinese BERT as our base student model. We compared our approach against two baselines: (1) a standard data distillation baseline, where BERT is fine-tuned directly on the LLM-generated labels; and (2) an embedding-alignment baseline (Previous Method) as discussed in Section 3.3.1, where the reasoning path is encoded using the BGE-M3 sentence embedding model (Chen et al., 2024) and aligned with BERT’s [CLS] representation. The alignment loss weight $\mu$ was set to 0.1. Regarding input construction, the standard student configuration follows the [CLS] query [SEP] service format with a maximum length of 64 tokens. For teacher configuration, we extended the maximum length to 150 by appending the reasoning path ([SEP] reason). All models were trained on 8 NVIDIA A100 GPUs using the PyTorch (Paszke et al., 2019) framework and the AdamW (Loshchilov and Hutter, 2017) optimizer. We used a constant learning rate of 5e-5, a per-GPU batch size of 512 (for an effective total batch size of 4,096), and trained all models for 5,000 steps. For CRSD method, the balancing hyperparameters for the teacher loss ($\gamma$) and the contrastive loss ($\delta$) were both set to 0.01.
@@ 5.1. LLM Results and Ablation Study
The offline experimental results for our domain-adapted reasoning LLM, presented in  Table 3, demonstrate a clear and steady performance improvement with each successive training stage. Our analysis begins with the SFT Baseline, which represents a model fine-tuned only on human-annotated reasoning data, without the benefit of our CPT stage.
The first significant enhancement comes from introducing domain-specific knowledge via Continued Pre-Training. The CPT+SFT model, which builds upon the CPT stage, shows a notable improvement across all metrics, with Macro F1 increasing from 0.6980 to 0.7032. This validates our initial hypothesis: equipping the LLM with foundational knowledge of our platform’s services and user behavior through a multi-task CPT phase is crucial before teaching it to perform complex reasoning. Next, we applied GRPO optimization, first using only the final answer’s correctness as a reward signal (CPT+SFT+GRPO($R_{label}$)). This led to another substantial performance jump, lifting Macro F1 to 0.7101. This result confirms that reinforcement learning is effective at refining the model’s policy. However, as discussed in our methods, relying on a sparse, outcome-only reward risks ”sparse reward” and ”reward hacking.”
@@
Table 3. Offline results for the reasoning LLM. The GRPO variants are applied on top of ‘CPT+SFT’, using either a simple label-based reward ($R_{label}$’) or our full weighted reward (‘Weighted’). The best results are highlighted in bold.
@@ 5.2. Distillation Results and Ablation Study
The results in  Table 4 illustrate the distinct contributions of each component. The + $\mathcal{L}_{\text{align}}$ experiment, which introduces only the in-batch contrastive alignment, already achieves a significant performance boost over the baselines. This is a noteworthy finding, as it demonstrates that forcing the student to mimic its reasoning-augmented state via contrastive learning is an effective strategy for knowledge internalization, even without direct supervision on how the teacher should use the reasoning path. Our final distilled model, CRSD Full, achieves a Macro F1 of 0.7076. This represents a remarkably successful distillation, as it retains approximately 98.6% of the teacher LLM’s performance in a model that is orders of magnitude smaller and faster.
@@
Furthermore, to verify that the performance gains are genuinely driven by the semantic content of the reasoning paths, rather than being an artifact of the model’s structure or longer input sequences, we conducted two additional sanity-check experiments. In the ”w/o Reasoning Path” setup, the teacher and student receive identical inputs. In the ”w/ Random Reasoning” setup, the teacher is fed an irrelevant reasoning path randomly sampled from the entire distillation dataset, simulating the effect of injecting uncorrelated noise into the reasoning process. The results are shown in Table 5.
Table 5. Ablation study on the importance of the reasoning path.
@@
The results compellingly demonstrate the importance of meaningful reasoning. In the w/o Reasoning Path setup, the contrastive loss becomes trivial, and the model’s performance regresses to be nearly identical to the standard Baseline. More strikingly, the w/ Random Reasoning experiment shows that providing noisy, irrelevant reasoning is actively detrimental, with performance dropping below the baseline.
@@ 5.3. Online Evaluation
To validate the practical value of our framework in a real-world business setting, we deployed the CRSD-enhanced lightweight model in Meituan’s search advertising system for a rigorous online A/B test. Over a one-week period, 10% of the live traffic was randomly allocated to the experimental group (our CRSD model), which was compared against a baseline group (the original 6-layer BERT model) comprising 30% of the traffic. The online experiment demonstrated comprehensive and statistically significant improvements across key business metrics, as summarized in Table 6.
@@
Ours +0.64% +1.73% +0.36% +0.10% +0.18%
The significant lift in Ad CTR and Ad CVR directly validates the model’s superior capability in query-ad relevance modeling. By more accurately matching user intent with platform services, our model effectively promoted user clicks and conversions, which in turn drove a steady increase in the platform’s Gross Transaction Volume (GTV). Furthermore, we conducted a manual analysis of the top-5 ad results for randomly sampled online queries. In a set of 3,000 samples, this analysis revealed a 0.8 percentage point reduction in the rate of relevance misjudgments.
@@ 6. Conclusion
In this paper, we proposed a two-stage framework to transfer the reasoning capabilities of Large Language Models (LLMs) to lightweight relevance models for industrial-scale search advertising. Our approach first constructs a domain-adapted reasoning LLM to generate high-quality labels and interpretable reasoning paths. Subsequently, our novel Contrastive Reasoning Self-Distillation (CRSD) method internalizes this reasoning knowledge into a compact BERT model, effectively overcoming the challenges of architectural heterogeneity. Comprehensive offline experiments demonstrate significant improvements in relevance prediction accuracy while maintaining high inference efficiency. Crucially, these offline gains were validated by online A/B testing in the Meituan search advertising system, which confirmed substantial improvements in key business metrics, user engagement, and a measurable reduction in relevance judgment errors. This work provides a practical and effective paradigm for integrating advanced LLM reasoning into production-grade systems, bridging the gap between performance, interpretability, and deployability in a real-world e-commerce search advertising context.
@@ Acknowledgements.
This research is supported by the National Natural Science Foundation of China (No. 62476127), the Meituan Research Fund (No. PO250624101698), the Natural Science Foundation of Jiangsu Province (No.BK20242039), the Basic Research Program of the Bureau of Science and Technology (ILF24001), the Scientific Research Starting Foundation of Nanjing University of Aeronautics and Astronautics (No.YQR21022).

### v3

@@ 4.1. Datasets and Evaluation
We employ three standard metrics to evaluate performance: (1) Accuracy, the proportion of correct predictions defined as $\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}(\hat{y}_{i}=y_{i})$, where $\mathbf{1}(\cdot)$ is the indicator function; (2) Macro F1, the unweighted average of F1-scores across all classes; and (3) Weighted F1, the average of class-wise F1-scores weighted by their respective support (sample counts).
@@ 4.2. Implementation Details
We select Qwen3-8B (Yang et al., 2025b) as the base model. During the CPT stage, training is conducted with a batch size of 32 and a learning rate of 1e-5, managed by a cosine scheduler with a warm-up ratio of 0.1. The AdamW optimizer is employed. The model is trained for 2 epochs with a maximum input length of 2,048 tokens, and the final loss converges to approximately 0.6. The hyperparameters for SFT are identical to those used in CPT, with the exception that the batch size is 16 and the number of epochs is set to 3. The reward model is trained using the same hyperparameters as the SFT stage. For the subsequent reinforcement learning (RL) stage, we employ the VERL framework (Sheng et al., 2025). During training, we sample 16 responses per prompt, using a prompt batch size of 64 (mini-batch size of 32) and the AdamW optimizer with a learning rate of 6e-7. To prevent GPU out-of-memory (OOM) errors, we deploy the reward model using vLLM (Kwon et al., 2023) and invoke it via API calls during the RL training process.
@@
For the distillation experiments, we use a 6-layer Chinese BERT as our base student model. We compare our approach against two baselines: (1) a standard data distillation baseline, where BERT is fine-tuned directly on the LLM-generated labels; and (2) an embedding-alignment baseline (Previous Method) as discussed in Section 3.3.1, where the reasoning path $r$ is encoded using the BGE-M3 sentence embedding model (Chen et al., 2024) and aligned with BERT’s [CLS] representation. The alignment loss weight $\mu$ is set to 0.1. Regarding input construction, the standard student configuration follows the [CLS] query [SEP] service format with a maximum length of 64 tokens. For the teacher configuration, we extend the maximum length to 150 by appending the reasoning path ([SEP] reason). All models are trained in a distributed data-parallel environment using the PyTorch (Paszke et al., 2019) framework and the AdamW (Loshchilov and Hutter, 2017) optimizer. We employ a constant learning rate of 5e-5 and an effective total batch size of 4,096, training for 5,000 steps. For the CRSD method, the balancing hyperparameters for the teacher loss ($\gamma$) and the contrastive loss ($\delta$) are both set to 0.01.
@@ 5.1. LLM Results and Ablation Study
We first evaluate the performance of general-purpose LLMs on our domain-specific task to verify the necessity of domain adaptation. As shown in the upper part of Table 3, state-of-the-art models like Qwen3-8B and DeepSeek-R1 (zero-shot and few-shot) exhibit significantly lower performance, with Accuracy hovering between 0.51 and 0.62. These models often struggle with platform-specific terminology and nuanced business logic. This substantial performance gap underscores that general-world knowledge alone is insufficient for high-precision e-commerce search, confirming the vital importance of our proposed domain-adaptive training pipeline. The bottom part of Table 3 presents the results of our reasoning LLM across different stages. It demonstrates a clear and steady performance improvement with each successive training step. Our analysis begins with the SFT Baseline, which represents a model fine-tuned only on human-annotated reasoning data without the benefit of the CPT stage.
The first significant enhancement comes from introducing domain-specific knowledge via Continued Pre-Training. The CPT+SFT model, which builds upon the CPT stage, shows a notable improvement across all metrics, with Macro F1 increasing from 0.6980 to 0.7032. This validates our initial hypothesis: equipping the LLM with foundational knowledge of our platform’s services and user behavior through a multi-task CPT phase is crucial before teaching it to perform complex reasoning. Next, we applied GRPO optimization, first using only the final answer’s correctness as a reward signal (CPT+SFT+GRPO($R_{label}$)). This led to another substantial performance jump, lifting Macro F1 to 0.7101. This result confirms that reinforcement learning is effective at refining the model’s policy. However, as discussed in our methods, relying on a sparse, outcome-only reward risks “sparse reward” and “reward hacking.”
@@
Table 3. Offline results for the reasoning LLM and general-purpose LLM baselines. The GRPO variants are applied on top of ‘CPT+SFT’, using either a simple label-based reward ($R_{label}$’) or our full weighted reward (‘Weighted’).
@@
Qwen3-8B (Zero-shot) 0.5480 0.3955 0.5748
Qwen3-8B (Few-shot) 0.5893 0.4375 0.6254
DeepSeek-R1 (Zero-shot) 0.5140 0.4067 0.5827
DeepSeek-R1 (Few-shot) 0.6197 0.4696 0.6638
@@ 5.2. Distillation Results and Ablation Study
The results in  Table 4 illustrate the distinct contributions of each component. The + $\mathcal{L}_{\text{align}}$ experiment, which introduces only the in-batch contrastive alignment, already achieves a significant performance boost over the baselines. This is a noteworthy finding, as it demonstrates that forcing the student to mimic its reasoning-augmented state via contrastive learning is an effective strategy for knowledge internalization, even without direct supervision on how the teacher should use the reasoning path. Our final distilled model, CRSD Full, achieves a Macro F1 of 0.7076. Notably, McNemar’s test results ($p<10^{-8}$) confirm that the improvement of CRSD over the baseline is statistically significant. This represents a remarkably successful distillation, as it retains approximately 98.6% of the teacher LLM’s performance in a model that is orders of magnitude smaller and faster.
@@
Furthermore, to verify that the performance gains are genuinely driven by the semantic content of the reasoning paths, rather than being an artifact of the model’s structure or longer input sequences, we conducted two additional sanity-check experiments. In the “w/o Reasoning Path” setup, the teacher and student receive identical inputs. In the “w/ Random Reasoning” setup, the teacher is fed an irrelevant reasoning path randomly sampled from the entire distillation dataset, simulating the effect of injecting uncorrelated noise into the reasoning process. The results are shown in Table 5.
Table 5. Ablation on the importance of reasoning path.
@@
The results compellingly demonstrate the importance of reasoning. In the w/o Reasoning Path setup, the contrastive loss becomes trivial, and the model’s performance regresses to be nearly identical to the standard Baseline. Strikingly, the w/ Random Reasoning experiment shows that providing noisy, irrelevant reasoning is actively detrimental, with performance dropping below the baseline.
@@ 5.3. Online Evaluation
To validate the practical value of our framework in a real-world business setting, we deployed the CRSD-enhanced lightweight model in Meituan’s search advertising system for a rigorous online A/B test. Over a two-week period, 10% of the live traffic was randomly allocated to the experimental group (our CRSD model), which was compared against a baseline group (the original 6-layer BERT model) comprising 30% of the traffic. The online experiment demonstrated comprehensive and statistically significant improvements across key business metrics, as summarized in Table 6.
@@
Ours +0.91% +1.06% +0.40% +0.11% +0.18%
The substantial improvements in Ad CTR and Ad CVR substantiate the model’s superior capability in query-ad relevance modeling. By aligning user intent more precisely with platform services, our approach effectively boosted user engagement, which consequently drove a steady increase in the platform’s Gross Transaction Volume (GTV). Furthermore, a manual expert review of the top 5 results from 5,521 randomly sampled queries revealed a 30.5 percentage point reduction in the bad case rate. Given these robust results, the model has been successfully deployed in the production environment.
@@ 6. Conclusion
In this paper, we proposed a two-stage framework to transfer the reasoning capabilities of Large Language Models (LLMs) to lightweight relevance models for industrial-scale search advertising. Our approach first constructs a domain-adapted reasoning LLM to generate high-quality labels and interpretable reasoning paths. Subsequently, our novel Contrastive Reasoning Self-Distillation (CRSD) method internalizes this reasoning knowledge into a compact BERT model, effectively overcoming the challenges of architectural heterogeneity. Comprehensive offline experiments demonstrate significant improvements in relevance prediction accuracy while maintaining high inference efficiency. Crucially, the offline gains were first validated through rigorous online A/B testing in the Meituan search advertising system, which showed substantial lifts in Ad CTR and Ad CVR, alongside a measurable reduction in the bad case rate. Following these robust results, our framework has been successfully deployed to production environment. This work provides a practical and effective paradigm for integrating advanced LLM reasoning into production-grade systems, bridging the gap between performance, interpretability, and deployability in a real-world e-commerce search advertising context.
@@ Acknowledgements.
This research is supported by the Meituan Research Fund, under Grant No. PO250624101698.